OpenStudy (anonymous):

How do I transform an absolute value equation into two linear equations? One of the problems is: Transform the absolute value equations into two linear equations. Then Solve |x-10|=17 I know what I would do if i just had to find x, but how do change it into 2 linear equations?

OpenStudy (freckles):

sorry my notifications aren't working what did you get for x for the other equation you were working?

OpenStudy (freckles):

like both x's

OpenStudy (anonymous):

for the other one I got that the extraneous solution was x=-2 and for this one like do I just write x= the two xs as the entire equation?

OpenStudy (freckles):

right but did you find the solution that was actually a solution for the other equation?

OpenStudy (freckles):

also again|x-10| can be written as (x-10) or -(x-10) so just like the other you will have two equations

OpenStudy (freckles):

x-10=17 or -(x-10)=17

OpenStudy (freckles):

whenever you have |f(x)| just say okay well that could either be f(x) or -(f(x))

OpenStudy (anonymous):

For the first one, the solution that was actually a solution would be 6=x, the other one? and for the second one, would the two linear equations be x-10=17 and -(x+10)=17 or do I need to solve and then use the 2 x answers as the linear equation?

OpenStudy (freckles):

ok good on the first one x=6 was a solution x=-2 was an extraneous solution it says to write as two separate linear equations which we did. And then it says solve. so you need then solve x-10=17 and solve also -(x+10)=17

OpenStudy (freckles):

so you have x-10=17 and x+10=-17 to solve

OpenStudy (freckles):

what do you get as solutions?

OpenStudy (anonymous):

I got x=27 and x=-7 I'm just confused about the whole linear equation thing

OpenStudy (freckles):

well x-10=17 and -(x+10)=17 are both linear equations (these equations aren't simplified) x=27 and x=-7 are both linear equations too (these linear equations are simplified)

OpenStudy (freckles):

wow and i wrote that one equation wrong twice -(x-10)=17*

OpenStudy (freckles):

x-10=-17 x=-17+10=-7 which is what you got

OpenStudy (anonymous):

So the answer is just x=27 and x=-7?

OpenStudy (freckles):

is this like where you enter into a computer software thing? or is this written work you turn into your teacher?

OpenStudy (freckles):

x=27 or x=-7 or the solutions and also linear equations

OpenStudy (anonymous):

It is written work. Why? I have a test coming up and I just want to make sure I'm understanding everything

OpenStudy (freckles):

if it is written work i think you to say well |x-10|=17 can be transformed into the two linear equations x-10=17 or x-10=-17. And then for the solving part you would x=27 or x=-7.

OpenStudy (anonymous):

Okay thanks!

OpenStudy (freckles):

np